OpenStudy (anonymous):

two cards are drawn without replacement from a standard 52 card deck. what is the probability of drawing two kings?

OpenStudy (anonymous):

i have 1/169 ; 1/221 ; 4/663 ; 3/676 as a choice of an answer

OpenStudy (anonymous):

The probability of drawing the first card is 4/52. Since the card has not been replaced, the probability of drawing the second card is 3/51. Thus the probability of drawing 2 kings without replacement is (4/52)(3/51) =1/221 = .00452.